HCl solution consists of 28% HCl by mass and has a density of 1.14g/ml. Find volume of the solution required to react with 1.87g Al as in the reaction:
2Al + 6HCl = 2AlCl3 + 3H2
1
Expert's answer
2020-12-10T07:45:12-0500

2Al + 6HCl = 2AlCl3 + 3H2

M (Al) = 26.98 g/mol

2 x 26.98 reacts

M (HCl) = 36.46 g/mol

6 x 36.46 g reacts

Given mass of Al: 1.87 g

The ratio is:

1.87 / 2x26.98 = x / 6x36.46

x = 7.58 g of HCl

The mass of the solution:

7.58 / 0.28 = 27.08 g

d = m/V

V = m/d

V = 27.08 / 1.14 = 23.75 mL